OpenStudy (anonymous):

:Please help, I really need help For the given quadratic equation convert into vertex form, find the vertex, and find the value for x = 6. Show your work. y = -2x2 + 2x +2

OpenStudy (anonymous):

A good start is factoring out a "2"

OpenStudy (anonymous):

can you take me through step by step

OpenStudy (anonymous):

@MathNinja_C you there? o:

OpenStudy (anonymous):

Can you look into your book and write the basic equation of a quadratic in vertex form

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

y = a(x – h)2 + k Does this look right

OpenStudy (anonymous):

y = a(x-h)^2 + k

OpenStudy (anonymous):

Sorry, not sure what happened to the first formula

OpenStudy (anonymous):

Yo need to get the original to match the formula of y = a(x-h)^2+k. I would start by moving the constant term of the other side and factoring out the two. 1) -2(x^2+x) = -2

OpenStudy (anonymous):

Can you que me into what might be the next step? Or are you confused

OpenStudy (anonymous):

Sorry, I did my factoring wrong -2(x^2-x) = -2

OpenStudy (perl):

$$ \Large y = -2x^2 + 2x +2 \\ \Large y = -2(x^2 +x) +2 \\ \Large y = -2(x^2 +x +\frac{1}{4} -\frac{1}{4} ) +2 \\ \Large y = -2((x+\frac{1}{2})^2 -\frac{1}{4} ) +2 \\ \Large y = -2(x+\frac{1}{2})^2 -2(-\frac{1}{4} ) +2 \\ \Large y = -2(x+\frac{1}{2})^2 +\frac{1}{2} +2 \\ \Large y = -2(x+\frac{1}{2})^2 +\frac{5}{2} \\ $$

OpenStudy (anonymous):

Or perl can just give you the answer